I had made a 5 year investment of N$ 50000, and earns N$ 120000,what was the prevailing interest rate than?
Answer by Theo(13342)

you invest 50,000 and get back 120,000 in 5 years.
if you compound annually, then the interest rate on your investment is:
50,000 * (1+x)^5 = 120,000
divide both sides of this equation by 50,000 to get:
(1+x)^5 = 12/5
take the 5th root of each side of this equation to get:
1+x = 1.191357898
subtract 1 from both sides of this equation to get:
x = .191357898
that's your annual interest rate.
50,000 * (1+.191357898)^5 equals 50,000 * (1.191357898)^5 which gets you 120,000 as it should.